Two resistor RA = 1.95ohms, and RB = 0.05 ohm are connected in parallel and take a total current of 50 amp. What is the current through each resistor.


1
Expert's answer
2022-01-07T14:33:01-0500

Given:

"R_A=1.95\:\Omega"

"R_B=0.05\:\Omega"

"I=50\:\rm A"


The parallel conection equations give

"I=I_A+I_B,\\ I_AR_A=I_BR_B."


In our case


"50=I_A+I_B,\\ 1.95I_A=0.05I_B."

We get

"I_B=39I_A"

"50=40I_A"

Finally

"I_A=50/40=1.25\:\rm A"

"I_B=39I_A=39*1.25\:\rm A=48.75\: A"